OpenStudy (anonymous):

Which of the following represents 6 x to the 3 fourths power in radical form?

563blackghost (563blackghost):

Can you put up the choices for this question?

OpenStudy (anonymous):

\[\sqrt[3]{6x}^{4}\] \[\sqrt[4]{6x}^{3}\] \[6\sqrt[3]{x}^{4}\] \[6\sqrt[4]{x}^{3}\]

563blackghost (563blackghost):

http://openstudy.com/study#/updates/53c4cc72e4b05c273e9d6608

OpenStudy (anonymous):

thanx

563blackghost (563blackghost):

Np ;) happy to help